OpenStudy (anonymous):

What is the gravitational field strength at a height h above the surface of the Earth? R is the radius of the earth

OpenStudy (anonymous):

G*M/(R+h)^2

OpenStudy (anonymous):

Gravitational field strength = Force / mass [ field strength is numerically equal to accelaration ] => E = [G.Me.m/r^2]/m = G.Me/r^2 Here, r = R+h => E = G.Me/(R+h)^2